### v1

@@ 4.3 Probabilistic Cell Problem and Overtaking Optimality
$J_{T}(x;R):=\mathbb{E}_{x}\bigg[\int_{0}^{T}\gamma c(X(s))\,ds+\sum_{k=1}^{\infty}I_{\{\tau_{k}\leq T\}}(p(X(\tau_{k}-)-X(\tau_{k}))-K)\bigg],$
@@
Such a problem is motivated by the so-called probabilistic cell problem studied in the context of LQG problem; see, for example, Bayraktar and Jian, (2025), Jian et al., (2024). Similar to the LQG setting, in which the probabilistic cell problem is related to the ergodic Bellman equation, we demonstrate that for the impulse control model studied in this paper, the solution to (4.20) is closely connected to the function $G$ defined in (4.10). In fact, recall that $(G,F^{*})$ solves the QVI (4.12) and that $F^{*}$ is the optimal value for the long-term average problem (1.4). Proposition 4.10 shows that if $a$ is an entrance boundary, then the value function $V$ of (4.20) is given by the impulse reward potential $G$, adjusted by its mean under the invariant measure resulting from the optimal $(w^{*},y^{*})$-policy.
@@
Assume Conditions 2.1, 2.5, 2.6, and 4.1 hold and that $a$ is an entrance boundary. Then the solution to problem (4.20) is the triplet $(F^{*},G(\cdot)-\langle G,\nu^{R^{*}}\rangle,R^{*})$, where $F^{*}=F(w^{*},y^{*})$ in which $(w^{*},y^{*})$ is the unique maximizing pair for the function $F$ of (3.3), $R^{*}=R^{(w^{*},y^{*})}$, $\nu^{R^{*}}$ is the invariant measure induced by $R^{*}$. Therefore the value function $V=G-\langle G,\nu^{R^{*}}\rangle$ is the deviation of the impulse reward potential from its mean under $\nu^{R^{*}}$.
@@
We have observed in Proposition 4.5 that the policy $R^{*}$ is optimal for (1.4) with optimal payoff $F^{*}$. We now show that the value function $V(\cdot)$ of (4.20) is given by $G(\cdot)-\langle G,\nu^{R^{*}}\rangle$ and that $R^{*}$ is also optimal for (4.20). In view of (3.3) and Proposition 4.9, for any $R\in{\mathcal{A}}_{\mathbb{T}}$ with $R^{(w,y)}\neq R^{*}$, we have $J(R^{(w,y)})=F(w,y)<F^{*}$. Therefore it follows that
@@
On the other hand, for the $R^{*}$-policy, denoting by $X$ the controlled process, using the same calculations as those in deriving (4.1), we obtain
$J_{T}(x;R^{*})-F^{*}T=G(x)-\lim_{n\to\infty}\mathbb{E}_{x}[G(X(T\wedge\beta_{n}))].$
Since $a$ is an entrance boundary, as observed in the proof of Lemma 4.4, the function $G$ is uniformly bounded from below. On the other hand, since the process $X$ is bounded above by $x_{0}\vee y^{*}$, the random variables $G(X(T\wedge\beta_{n})),n\in\mathbb{N}$ are uniformly bounded. Hence $\lim_{n\to\infty}\mathbb{E}_{x}[G(X(T\wedge\beta_{n}))]=\mathbb{E}_{x}[G(X(T))]$. Furthermore, since $X$ possesses the invariant measure $\nu^{R^{*}}$, we have $\lim_{T\to\infty}\mathbb{E}_{x}[G(X(T))]=\langle G,\nu^{R^{*}}\rangle$. Using this observation in the above displayed equation yields
$\liminf_{T\to\infty}[J_{T}(x;R^{*})-F^{*}T]=G(x)-\langle G,\nu^{R^{*}}\rangle.$ (4.22)
A combination of (4.21) and (4.22) yields $V(\cdot)=G(\cdot)-\langle G,\nu^{R^{*}}\rangle$ and that $R^{*}$ is an optimal policy for (4.20). This completes the proof. ∎
@@
$\limsup_{T\to\infty}[J_{T}(x;R)-F^{*}T]$ $\leq G(x)-\liminf_{T\to\infty}\lim_{n\to\infty}\mathbb{E}_{x}[G(X(T\wedge\beta_{n}))]$
$\leq G(x)-\liminf_{T\to\infty}\mathbb{E}_{x}[G(X(T))]$
$\leq G(x)-\langle G,\nu^{R}\rangle,$
where the second inequality follows from Fatou’s lemma and the assumption that $a$ is an entrance boundary, which ensures that $G$ is bounded from below; which together with the continuity of $G$ and the Portmanteau theorem imply the last inequality. Therefore it follows that
$V(x)\leq G(x)-\inf_{R\in{\mathcal{A}}_{\mathrm{S}}:J(R)=F^{*}}\langle G,\nu^{R}\rangle.$ (4.23)
@@
We now turn to another perspective on impulse control over an infinite horizon: overtaking optimality. This important concept in infinite-horizon optimization was first introduced in Ramsey, (1928), with a weaker formulation later provided in von Weizsäcker, (1965) and others. Following the terminology in Huang et al., (2021), we say that an admissible policy $\widehat{R}\in{\mathcal{A}}$ is overtaking optimal starting from $x\in{\mathcal{I}}$ if for each $R\in{\mathcal{A}}$, we have
@@ 5 The Singular Control Problem
$\lim_{t\to\infty}t^{-1}\mathbb{E}_{x_{0}}\bigg[\int_{0}^{t}\gamma c(X^{Z_{x}}(s))ds\bigg]=\int_{a}^{b}\gamma c(u)\pi_{x}(u)\,du.$ (5.4)
@@
$\mathbb{E}_{x_{0}}[X^{Z_{x}}(t)]$ $=x_{0}+\mathbb{E}_{x_{0}}\bigg[\int_{0}^{t}\mu(X^{Z_{x}}(s))ds-(x-x_{0})^{+}-L_{x}(t)\bigg].$
@@
$0$ $=\lim_{t\to\infty}t^{-1}\mathbb{E}_{x_{0}}\bigg[\int_{0}^{t}\mu(X^{Z_{x}}(s))ds\bigg]-\lim_{t\to\infty}t^{-1}\mathbb{E}_{x_{0}}[L_{x}(t)]$
@@
$\mathbb{E}_{x_{0}}[U(X(T\wedge\beta_{n})])$ $=U(x_{0})+\mathbb{E}_{x_{0}}\bigg[\int_{0}^{T\wedge\beta_{n}}AU(X(s))ds-\int_{0}^{T\wedge\beta_{n}}U^{\prime}(X(s-))\,dZ({s})^{c}$
$\hskip 18.49988pt+\sum_{0\leq s\leq T\wedge\beta_{n}}\left[U(X(s))-U(X(s-))\right]\bigg].$ (5.9)
@@
$\mathbb{E}_{x_{0}}[U(X(T\wedge\beta_{n}))]$ $\leq U(x_{0})-\mathbb{E}_{x_{0}}\bigg[\int_{0}^{T\wedge\beta_{n}}\gamma c(X(s))ds+h(\widehat{y})(T\wedge\beta_{n})-pZ(T\wedge\beta_{n})\bigg].$ (5.11)
Using an argument analogous to that in the proof of Lemma 4.4, it follows from (5.1) that $\liminf_{t\to\infty}\liminf_{n\to\infty}T^{-1}\mathbb{E}_{x}[U(X(T\wedge\beta_{n}))]\geq 0$. Therefore, by rearranging terms of (5.11), dividing both sides by $T$, and then sending $T\to\infty$, we obtain from the monotone convergence theorem that
@@ 6 Sensitivity Results
$F^{*}(K):=F(K;w_{K}^{*},y_{K}^{*})=\bigg(\frac{pB\mathfrak{i}}{B\xi}+\frac{\gamma Bg}{B\xi}-\frac{K}{B\xi}\bigg)(w_{K}^{*},y_{K}^{*})$
@@
a. Suppose $w^{*}_{p}\geq\widetilde{x}_{\mu}$. Since $h_{\mu}$ is strictly decreasing on $(\widetilde{x}_{\mu},b)$ and $\theta_{p}>w^{*}_{p}$, it follows that $h_{\mu}(\theta_{p})-h_{\mu}(w^{*}_{p})<0$ and hence that $\dot{w}^{*}_{p}<0$.
b. Suppose $w^{*}_{p}<\widetilde{x}_{\mu}$. This analysis compares $h_{\mu}$ on $w^{*}_{p},b)$ to the function $y\mapsto\mathfrak{z}(w^{*}_{p},y)=\frac{B\mathfrak{i}}{B\xi}(w^{*}_{p},y)$ for $y>w^{*}_{p}$. Several subcases need to be examined.
First, applying the extended mean value theorem, for each $y>w^{*}_{p}$, there is some $\theta_{y}$ with $w^{*}_{p}<\theta_{y}<y$ for which $\mathfrak{z}(w^{*}_{p},y)=\frac{1}{\xi^{\prime}(\theta_{y})}=h_{\mu}(\theta_{y})$. Extending $\mathfrak{z}(w^{*}_{p},y)$ by continuity to $y=w^{*}_{p}$, we have $\mathfrak{z}(w^{*}_{p},w^{*}_{p})=h_{\mu}(w^{*}_{p})$. In addition, on the interval $(w^{*}_{p},\widetilde{x}_{\mu})$, $h_{\mu}$ is increasing so since $w^{*}_{p}<\theta_{y}$, it follows that $h_{\mu}(w^{*}_{p})<\mathfrak{z}(w^{*}_{p},y)$. If $\theta_{p}\in(w^{*}_{p},\widetilde{x}_{\mu})$, then

### v2

@@ 4.3 Probabilistic Cell Problem and Overtaking Optimality
$J_{T}(x;R):=\mathbb{E}_{x}\left[\int_{0}^{T}\gamma c(X(s))\,ds+\sum_{k=1}^{\infty}I_{\{\tau_{k}\leq T\}}(p(X(\tau_{k}-)-X(\tau_{k}))-K)\right],$
@@
Such a problem is motivated by similar problems arising in Statistics and the so-called probabilistic cell problem studied in the context of the LQG problem; see, for example, Bayraktar and Jian, (2025), Jian et al., (2024). Similar to the LQG setting, in which the probabilistic cell problem is related to the ergodic Bellman equation, we demonstrate that for the impulse control model studied in this paper, the solution to (4.20) is closely connected to the function $G$ defined in (4.10). In fact, recall that $(G,F^{*})$ solves the QVI (4.12) and that $F^{*}$ is the optimal value for the long-term average problem (1.4). Proposition 4.10 shows that if $a$ is an entrance boundary, then the value function $V$ of (4.20) is given by the impulse reward potential $G$, adjusted by its mean under the invariant measure resulting from the optimal $(w^{*},y^{*})$-policy; further adjustment is needed when the initial state $x$ is above the threshold $y^{*}$.
@@
Assume the conditions of Proposition 4.9, let $(w^{*},y^{*})$ be the unique optimizer of the function $F$ of (3.3) and, as before, denote $F(w^{*},y^{*})=F^{*}$. We now define piecewisely the function $\widetilde{G}$ as follows:
$\widetilde{G}(x)=\begin{cases}F^{*}\xi(x)-\gamma g(x),&x<y^{*},\\ p(x-w^{*})-K+G(w^{*}),&x\geq y^{*}.\end{cases}$ (4.21)
Notice that $\widetilde{G}(x)=G(x)$ for $x\leq y^{*}$, where the function $G$ is defined in (4.10), but then $\widetilde{G}$ grows linearly for $x\geq\ y^{*}$. Hence we can regard $\widetilde{G}$ as a modified impulse reward functional.
@@
Assume Conditions 2.1, 2.5, 2.6, and 4.1 hold and that $a$ is an entrance boundary. Let $\widetilde{G}$ be defined by (4.21) with $(w^{*},y^{*})$ being the unique optimizer of $F$. Let $R^{*}$ denote the optimal $(w^{*},y^{*})$-policy and $\nu^{R^{*}}$ be the invariant measure induced by $R^{*}$. Then the solution to problem (4.20) is the triplet
$(F^{*},\widetilde{G}(\cdot)-\langle\widetilde{G},\nu^{R^{*}}\rangle,R^{*}).$
Therefore the value function $V=\widetilde{G}-\langle\widetilde{G},\nu^{R^{*}}\rangle$ is the deviation of the modified impulse reward potential from its mean under $\nu^{R^{*}}$.
@@
We have observed in Proposition 4.5 that the policy $R^{*}$ is optimal for (1.4) with optimal payoff $F^{*}$. We now show that the value function $V(\cdot)$ of (4.20) is given by $\widetilde{G}(\cdot)-\langle\widetilde{G},\nu^{R^{*}}\rangle$ and that $R^{*}$ is also optimal for (4.20). In view of (3.3) and Proposition 4.9, for any $R\in{\mathcal{A}}_{\mathbb{T}}$ with $R^{(w,y)}\neq R^{*}$, we have $J(R^{(w,y)})=F(w,y)<F^{*}$. Therefore it follows that
@@
On the other hand, for the $R^{*}$-policy, denoting by $X$ the controlled process, the same calculations as those in deriving in (4.1) give
$G(x)$ $+F^{*}T-\liminf_{n\to\infty}\mathbb{E}_{x}[G(X(T\wedge\beta_{n}))]$
$=\mathbb{E}_{x}\left[\int_{0}^{T}\gamma c(X(s))ds+\sum_{k=1}^{\infty}I_{\{\tau_{k}\leq T\}}[G(X(\tau_{k}-))-G(X(\tau_{k}))]\right].$ (4.23)
We proceed in two cases.
  • (i)
    If $x<y^{*}$, then under the policy $R^{*}$, $X(\tau_{k}-)=y^{*}$ and $X(\tau_{k})=w^{*}$ for all $k\in\mathbb{N}$. In addition, the last equaltion of (4.11) implies that
    $G(X(\tau_{k}-))-G(X(\tau_{k}))$ $=G(y^{*})-G(w^{*})=p(y^{*}-w^{*})-K$
    $=p(X(\tau_{k}-)-X(\tau_{k}))-K,\qquad k\in\mathbb{N}.$ (4.24)
    Plugging ((i)) into (4.3) yields
    $G$ $(x)+F^{*}T-\lim_{n\to\infty}\mathbb{E}_{x}[G(X(T\wedge\beta_{n}))]$
    $=\mathbb{E}_{x}\left[\int_{0}^{T}\gamma c(X(s))ds+\sum_{k=1}^{\infty}I_{\{\tau_{k}\leq T\}}[p(X(\tau_{k}-)-X(\tau_{k}))-K]\right].$
    Rearrange the above displayed equation to obtain
    $J_{T}(x;R^{*})-F^{*}T$ $=G(x)-\lim_{n\to\infty}\mathbb{E}_{x}[G(X(T\wedge\beta_{n}))]=\widetilde{G}(x)-\lim_{n\to\infty}\mathbb{E}_{x}[G(X(T\wedge\beta_{n}))],$ (4.25)
    where the last equality holds because $\widetilde{G}(x)=G(x)$ for all $x\leq y^{*}$.
  • (ii)
    If $x\geq y^{*}$, then $X$ jumps to $w^{*}$ at time $\tau_{1}=0$ and subsequently behaves as if it started from $w^{*}$. Therefore,
    $G$ $(X(\tau_{1}-))-G(X(\tau_{1}))$
    $=G(x)-G(w^{*})=p(x-w^{*})-K+G(x)-G(w^{*})-[p(x-w^{*})-K]$
    $=p(X(\tau_{1}-)-X(\tau_{1}))-K+G(x)-G(w^{*})-[p(x-w^{*})-K];$ (4.26)
    ((i)) still holds for all $k\geq 2$. Using ((ii)) and ((i)) in (4.3), we derive
    $G$ $(x)+F^{*}T-\liminf_{n\to\infty}\mathbb{E}_{x}[G(X(T\wedge\beta_{n}))]$
    $=\mathbb{E}_{x}\left[\int_{0}^{T}\gamma c(X(s))ds+\sum_{k=1}^{\infty}I_{\{\tau_{k}\leq T\}}[p(X(\tau_{k}-)-X(\tau_{k}))-K]\right]$
    $\qquad+G(x)-G(w^{*})-[p(x-w^{*})-K].$
    Using the definition of $\widetilde{G}$ given in (4.21), a rearrangement of the above displayed equation leads to
    $J_{T}(x;R^{*})-F^{*}T$ $=G(w^{*})+[p(x-w^{*})-K]-\lim_{n\to\infty}\mathbb{E}_{x}[G(X(T\wedge\beta_{n}))]$
    $=\widetilde{G}(x)-\lim_{n\to\infty}\mathbb{E}_{x}[G(X(T\wedge\beta_{n}))];$
    again establishing (4.25)
Since $a$ is an entrance boundary, as observed in the proof of Lemma 4.4, the function $G$ is uniformly bounded from below. On the other hand, since the process $X$ is bounded above by $x_{0}\vee y^{*}$, the random variables $G(X(T\wedge\beta_{n})),n\in\mathbb{N}$ are uniformly bounded. Hence $\lim_{n\to\infty}\mathbb{E}_{x}[G(X(T\wedge\beta_{n}))]=\mathbb{E}_{x}[G(X(T))]$. Furthermore, since $X$ possesses the invariant measure $\nu^{R^{*}}$, we have $\lim_{T\to\infty}\mathbb{E}_{x}[G(X(T))]=\langle G,\nu^{R^{*}}\rangle$. Since $\nu^{R^{*}}$ has support on $a,y^{*}]$ and $G$ and $\widetilde{G}$ coinside on this interval, we have $\langle G,\nu^{R^{*}}\rangle=\langle\widetilde{G},\nu^{R^{*}}\rangle$. Therefore taking the limit as $T\to\infty$ in (4.25) yields
$\liminf_{T\to\infty}[J_{T}(x;R^{*})-F^{*}T]=\widetilde{G}(x)-\langle G,\nu^{R^{*}}\rangle=\widetilde{G}(x)-\langle\widetilde{G},\nu^{R^{*}}\rangle.$ (4.27)
A combination of (4.22) and (4.27) yields $V(\cdot)=\widetilde{G}(\cdot)-\langle\widetilde{G},\nu^{R^{*}}\rangle$ and that $R^{*}$ is an optimal policy for (4.20). This completes the proof. ∎
Remark 4.11.
As a side remark, we note that the function $\widetilde{G}$ defined in (4.21) satisfies $\widetilde{G}\in C^{1}({\mathcal{I}})\cap C^{2}({\mathcal{I}}\backslash\{y^{*}\})$. Moreover, $\widetilde{G}$ and $F^{*}$ satisfy the system (4.11) and the QVI (4.12). In fact, in view of the smooth pasting method for ergodic singular and impulse control problems, c.f. Helmes et al., (2017), Jack and Zervos, (2006), Kunwai et al., (2022), Liang et al., (2025), we may regard $(\widetilde{G},F^{*})$ as the usual solution to (4.12).
@@
$\limsup_{T\to\infty}[J_{T}(x;R)-F^{*}T]$ $\leq\widetilde{G}(x)-\liminf_{T\to\infty}\lim_{n\to\infty}\mathbb{E}_{x}[\widetilde{G}(X(T\wedge\beta_{n}))]$
$\leq\widetilde{G}(x)-\liminf_{T\to\infty}\mathbb{E}_{x}[\widetilde{G}(X(T))]$
$\leq\widetilde{G}(x)-\langle\widetilde{G},\nu^{R}\rangle,$
where the second inequality follows from Fatou’s lemma and the assumption that $a$ is an entrance boundary, which ensures that $\widetilde{G}$ is bounded from below, and which together with the continuity of $\widetilde{G}$ and the Portmanteau theorem imply the last inequality. Therefore it follows that
$V(x)\leq\widetilde{G}(x)-\inf_{R\in{\mathcal{A}}_{\mathrm{S}}:J(R)=F^{*}}\langle\widetilde{G},\nu^{R}\rangle.$ (4.28)
@@
We now turn to another perspective on impulse control over an infinite horizon: overtaking optimality. This important concept in infinite-horizon optimization was first introduced in Ramsey, (1928) and has applications in economics, operation research, and control theory. Weaker formulations were subsequently developed in Brock and Haurie, (1976), von Weizsäcker, (1965), Gale, (1967) and many other works; see also Dockner et al., (2000), Carlson et al., (1991) for extensive surveys and comprehensive lists of references on overtaking optimality. Following the terminology in Huang et al., (2021), we say that an admissible policy $\widehat{R}\in{\mathcal{A}}$ is overtaking optimal starting from $x\in{\mathcal{I}}$ if for each $R\in{\mathcal{A}}$, we have
@@ 5 The Singular Control Problem
$\lim_{t\to\infty}t^{-1}\mathbb{E}_{x_{0}}\left[\int_{0}^{t}\gamma c(X^{Z_{x}}(s))ds\right]=\int_{a}^{b}\gamma c(u)\pi_{x}(u)\,du.$ (5.4)
@@
$\mathbb{E}_{x_{0}}[X^{Z_{x}}(t)]$ $=x_{0}+\mathbb{E}_{x_{0}}\left[\int_{0}^{t}\mu(X^{Z_{x}}(s))ds-(x-x_{0})^{+}-L_{x}(t)\right].$
@@
$0$ $=\lim_{t\to\infty}t^{-1}\mathbb{E}_{x_{0}}\left[\int_{0}^{t}\mu(X^{Z_{x}}(s))ds\right]-\lim_{t\to\infty}t^{-1}\mathbb{E}_{x_{0}}[L_{x}(t)]$
@@
$\mathbb{E}_{x_{0}}[U(X(T\wedge\beta_{n})])$ $=U(x_{0})+\mathbb{E}_{x_{0}}\left[\int_{0}^{T\wedge\beta_{n}}AU(X(s))ds-\int_{0}^{T\wedge\beta_{n}}U^{\prime}(X(s-))\,dZ({s})^{c}\right$.
$\hskip 18.49988pt\left.+\sum_{0\leq s\leq T\wedge\beta_{n}}\left[U(X(s))-U(X(s-))\right]\right].$ (5.9)
@@
$\mathbb{E}_{x_{0}}[U(X(T\wedge\beta_{n}))]$ $\leq U(x_{0})-\mathbb{E}_{x_{0}}\left[\int_{0}^{T\wedge\beta_{n}}\gamma c(X(s))ds+h(\widehat{y})(T\wedge\beta_{n})-pZ(T\wedge\beta_{n})\right].$ (5.11)
Using an argument analogous to that in the proof of Lemma 4.4, it follows from (5.1) that
$\liminf_{t\to\infty}\liminf_{n\to\infty}T^{-1}\mathbb{E}_{x}[U(X(T\wedge\beta_{n}))]\geq 0.$
Therefore, by rearranging terms of (5.11), dividing both sides by $T$, and then sending $T\to\infty$, we obtain from the monotone convergence theorem that
@@ 6 Sensitivity Results
$F^{*}(K):=F(K;w_{K}^{*},y_{K}^{*})=\left(\frac{pB\mathfrak{i}}{B\xi}+\frac{\gamma Bg}{B\xi}-\frac{K}{B\xi}\right)(w_{K}^{*},y_{K}^{*})$
@@
(a) Suppose $w^{*}_{p}\geq\widetilde{x}_{\mu}$. Since $h_{\mu}$ is strictly decreasing on $(\widetilde{x}_{\mu},b)$ and $\theta_{p}>w^{*}_{p}$, it follows that $h_{\mu}(\theta_{p})-h_{\mu}(w^{*}_{p})<0$ and hence that $\dot{w}^{*}_{p}<0$.
(b) Suppose $w^{*}_{p}<\widetilde{x}_{\mu}$. This analysis compares $h_{\mu}$ on $w^{*}_{p},b)$ to the function $y\mapsto\mathfrak{z}(w^{*}_{p},y)=\frac{B\mathfrak{i}}{B\xi}(w^{*}_{p},y)$ for $y>w^{*}_{p}$. Several subcases need to be examined. First, applying the extended mean value theorem, for each $y>w^{*}_{p}$, there is some $\theta_{y}$ with $w^{*}_{p}<\theta_{y}<y$ for which $\mathfrak{z}(w^{*}_{p},y)=\frac{1}{\xi^{\prime}(\theta_{y})}=h_{\mu}(\theta_{y})$. Extending $\mathfrak{z}(w^{*}_{p},y)$ by continuity to $y=w^{*}_{p}$, we have $\mathfrak{z}(w^{*}_{p},w^{*}_{p})=h_{\mu}(w^{*}_{p})$. In addition, on the interval $(w^{*}_{p},\widetilde{x}_{\mu})$, $h_{\mu}$ is increasing so since $w^{*}_{p}<\theta_{y}$, it follows that $h_{\mu}(w^{*}_{p})<\mathfrak{z}(w^{*}_{p},y)$. If $\theta_{p}\in(w^{*}_{p},\widetilde{x}_{\mu})$, then